A solid sphere of mass 100 kg and radius 10 m moving in a space becomes a circular disc of radius 20 m in one hour. Then the rate of change of moment of inertia in the process is
1
Expert's answer
2019-08-29T09:53:44-0400
"I=\frac{2}{5}mr^2"

"I'=\frac{1}{2}mR^2"

The rate of change of moment of inertia in the process is


"\frac{I'-I}{t}=\frac{\frac{1}{2}mR^2-\frac{2}{5}mr^2}{t}=m\frac{\frac{1}{2}R^2-\frac{2}{5}r^2}{t}"

"\frac{I'-I}{t}=100\frac{\frac{1}{2}20^2-\frac{2}{5}10^2}{3600}=4.44\frac{kg\cdot m^2}{s}"